# DualOptim: Enhancing Efficacy and Stability in Machine Unlearning with Dual Optimizers

### v1

@@ 3.4 Dual Optimizers for Machine Unlearning
2:  $\mathrm{Optim}$ is the same optimizer as in pretraining with step size $\eta$. $\mathrm{Optim}_{f}$ is $\mathrm{Adam}(\theta,\eta_{f})$$\mathrm{Optim}_{r}$ is the same optimizer as in pretraining with step size $\eta_{r}$.
@@ 4.1 Random Subset Unlearning in Image Classification
Table 1: Performance summary of MU methods for image classification. Experiments are conducted on (a) $10\%$ random subset of CIFAR-10 using ResNet-18 and (b) $10\%$ random subset of TinyImageNet using Swin-T. All results are presented as mean and standard deviation across $5$ trials with different random forget data. Performance gaps with RT are indicated in blue. The average gap (Gap) and average standard deviation (Std) metrics are calculated by the average of the gaps and standard deviation measured in FA, RA, TA, and MIA, respectively. All the numbers are in percentage.
@@
We start with random subset unlearning tasks in image classification, including using ResNet-18 [44] on CIFAR-10 [45] and Swin Transformer-Tiny (Swin-T) [46] on TinyImageNet [47]. Additional results on CIFAR-100 [45] and SVHN [48] are included in Appendix C.2. Consistent with previous work [21, 24, 6, 7], we regard Retrain (RT) as the gold standard of MU. Since our proposed DualOptim is plug-and-play, we apply it to SCRUB [8], SalUn [6], and SFRon [7] to validate its efficacy and stability. Additionally, we include three simple baselines, namely Fine-tune (FT) [49], Gradient Ascent (GA) [21] and Random Label (RL) [22], for reference as well. Note that the hyperparameters of all evaluated methods are tuned to their optimal values by sophisticated search, and we defer the implementation details to Appendix B. Besides the metrics mentioned in Sec. 3.1, which includes FA, RA, TA and MIA, we follow the evaluation criteria in [6, 7] to report the average gap (Gap) and average standard deviation (Std) to indicate the overall performance and stability, respectively. They are the average gap with RT and the average standard deviation among the results in FA, RA, TA, and MIA, respectively.
@@ 4.2 Class-wise Unlearning in Image Generation
Table 2: Class-wise unlearning performance on CIFAR-10 with DDPM and ImageNet with DiT. The best unlearning performance for each forgetting class is highlighted in bold for FA (in %) and FID. Note that the results of SA, SalUn and SFRon are those reported in [7].
Method CIFAR-10 Class-wise Unlearning ImageNet Class-wise Unlearning
Automobile Cat Dog Horse Truck Cockatoo Golden Retriever White Wolf Arctic Fox Otter
FA $\downarrow$ FID $\downarrow$ FA $\downarrow$ FID $\downarrow$ FA $\downarrow$ FID $\downarrow$ FA $\downarrow$ FID $\downarrow$ FA $\downarrow$ FID $\downarrow$ FA $\downarrow$ FID $\downarrow$ FA $\downarrow$ FID $\downarrow$ FA $\downarrow$ FID $\downarrow$ FA $\downarrow$ FID $\downarrow$ FA $\downarrow$ FID $\downarrow$
SA 0.00 23.56 14.20 21.34 8.60 21.19 0.00 21.13 0.00 29.04 0.00 348.75 0.00 298.97 0.00 45.89 0.00 393.91 29.8 321.21
SalUn 0.20 21.23 1.40 20.29 0.00 20.18 0.60 20.70 0.80 20.45 91.21 18.47 46.09 25.28 0.00 15.16 45.90 408.07 87.5 19.69
SFRon 0.00 20.70 7.40 18.44 0.20 18.89 0.00 19.93 0.00 20.61 0.00 13.59 0.00 17.76 0.00 23.28 0.00 16.12 0.00 16.43
+DO 0.20 19.72 1.00 19.36 0.00 18.58 0.00 18.91 0.00 17.26 0.00 17.46 0.00 14.63 0.00 14.72 0.00 14.91 0.00 14.55
Besides classification, we further evaluate our proposed DualOptim in class-wise unlearning tasks in image generation. Following the settings of [7], we conduct experiments using conditional DDPM [26] with the U-Net [50] on CIFAR-10 and the latent diffusion model [28] with Diffusion Transformer (DiT) [29] on ImageNet [47]. We apply our proposed DualOptim to SFRon [7], which exhibit the state-of-the-art unlearning performance in image generation. We also include SA [35] and SalUn [6] as baselines for comparison. Note that we do not apply DualOptim to SA and SalUn due to their joint updating scheme and unstable performance on ImageNet. Nonetheless, we still present the results of SalUn+DO on CIFAR-10 in Appendix C.3. The implementation details can be found in Appendix B. As for the evaluation metrics, we adopt the accuracy of the unlearned model’s generated images on forgetting classes (FA) by a pre-trained classifier to indicate the forgetting efficacy and Fréchet Inception Distance (FID) [51] to assess the generation capability on the retained classes.
@@ 4.3 Random Subset Unlearning in Large Language Models
Method Phi-1.5 LLaMA 2
forget 1% data forget 5% data forget 10% data forget 1% data forget 5% data forget 10% data
@@
GA+GD 0.4934 0.4493 0.4714 0.4360 0.5084 0.4722 0.4471 0.5246 0.4859 0.6696 0.5908 0.6302 0.0000 0.8772 0.4386 0.5592 0.9346 0.7469
ME+GD 0.4944 0.3938 0.4441 0.4559 0.4480 0.4520 0.4594 0.4564 0.4579 0.7271 0.9204 0.8237 0.7472 0.9313 0.8392 0.7357 0.9489 0.8423
+DO 0.4866 0.6913 0.5889 0.4676 0.8200 0.6438 0.5009 0.7732 0.6370 0.7425 0.9612 0.8519 0.7316 0.9602 0.8459 0.7315 0.9625 0.8470
DPO+GD 0.2410 0.6831 0.4621 0.4105 0.6334 0.5219 0.3517 0.6302 0.4910 0.7564 0.5335 0.6450 0.0000 0.8243 0.4122 0.0000 0.8041 0.4021
IDK+AP 0.4403 0.5723 0.5063 0.4800 0.5112 0.4956 0.4614 0.6003 0.5308 0.7580 0.7625 0.7603 0.7529 0.7479 0.7504 0.7471 0.7433 0.7452
+DO 0.4221 0.7037 0.5629 0.4633 0.6974 0.5804 0.4422 0.7193 0.5807 0.7412 0.8075 0.7743 0.7354 0.7958 0.7656 0.7362 0.7855 0.7609
Following the success of MU in vision tasks, there is a rising interest of applying it to remove private or harmful information from large language models (LLMs). We conduct the experiments on Phi-1.5 [52] and LLaMA 2 [37], both are fine-tuned on TOFU dataset [2] with 1.3B and 7B parameters, respectively. We include two untargeted unlearning methods (GA+GD and ME+GD) and two targeted unlearning methods (DPO+GD and IDK+AP) [9] as the baselines. We employ DualOptim (DO) in ME+GD and IDK+AP, which perform the best in [9], to validate the effectiveness of our method. Consistent with [2, 9], we consider three levels of unlearning tasks: to forget 1%, 5%, and 10% of the constructed data. We follow [9] and adopt the improved Model Capability (MC) 333 Model Capability (MC) is referred to Model Utility (MU) in [9], which conflicts with the abbreviation of Machine Unlearning (MU) in this paper. and Forger Efficacy (FE) as the evaluation metrics.
@@
It is important to note that the forgetting and retaining objectives are jointly optimized in the baselines listed in Table 4, whereas our method alternates between optimizing these two objectives. To ensure a fair comparison, we also evaluate unlearning performance by alternately optimizing the forgetting and retaining objectives using a shared optimizer. The results in Table 14 of Appendix C.4 demonstrate that DualOptim improves performance for both the joint update scheme and the alternate update scheme.
Although DualOptim exhibits effectiveness in MU for LLMs, the decoupled momentum introduces significant computational and memory overhead because of large amount of parameters in LLMs. To investigate our method’s effectiveness under limited budgets, we apply DualOptim to the parameter-efficient fine-tuning techniques, such as LoRA [53]. As indicated in Table 15 of Appendix C.4, DualOptim achieves a minimal compromise in unlearning performance by using LoRA, while significantly reducing memory consumption.
@@ 4.4 Ablation Studies
We compare various combinations of the Adam and SGD optimizers for the forget loss and the retain loss with their standalone counterparts. Besides adaptive learning rate optimizers, we investigate the combinations of a gradient sign-based optimizer, Lion [54], and SGD as well. The results, as shown in Table 5, indicate that the configuration of Adam (F) + SGD (R) offers the optimal balance between performance and stability. This finding underscores the importance of decoupled momentum and adaptive learning rate for the forgetting objective. Moreover, the results presented in Table 17 and Table 18 in Appendix C.5 indicate that the best optimizer for retaining depends on the choice of optimizer during pretraining and the specific MU algorithms employed. It should be pointed out that although a shared SGD demonstrates a competitive average performance, it suffers from high standard deviation indicating instability. Conversely, while Lion exhibits low variability, it does not achieve optimal performance. These insights further emphasize the efficacy of our method in enhancing performance and stability in MU tasks.
Table 5: Ablation study on different combinations of DualOptim. Results are based on $10\%$ random subset unlearning task on CIFAR-10 using ResNet-18 pre-trained by SGD. SFRon is the adopted MU algorithm. (F) and (R) denotes that the optimizer is used to minimize the forget and retain losses, respectively.
@@ Appendix A Proofs
[…]
@@
[…]
@@
$\mathrm{(Shared~{}Momentum)}$$\begin{cases}{\bm{m}}_{f,t}^{S}=\alpha{\bm{m}}_{r,t-1}^{S}+\widehat{{\bm{g}}}^{S}_{f,t}=\sum_{i=0}^{t}\alpha^{2(t-i)}\widehat{{\bm{g}}}^{S}_{f,i}+\sum_{i=0}^{t-1}\alpha^{2(t-i)-1}\widehat{{\bm{g}}}^{S}_{r,i}\\ {\bm{m}}_{r,t}^{S}=\alpha{\bm{m}}_{f,t}^{S}+\widehat{{\bm{g}}}^{S}_{r,t}=\sum_{i=0}^{t}\alpha^{2(t-i)+1}\widehat{{\bm{g}}}^{S}_{f,i}+\sum_{i=0}^{t}\alpha^{2(t-i)}\widehat{{\bm{g}}}^{S}_{r,i}\end{cases}$ (10)
$\mathrm{(Decoupled~{}Momentum)}$$\begin{cases}{\bm{m}}_{f,t}^{D}=\alpha{\bm{m}}_{f,t-1}^{D}+\widehat{{\bm{g}}}^{D}_{f,t}=\sum_{i=0}^{t}\alpha^{t-i}\widehat{{\bm{g}}}^{D}_{f,i}\\ {\bm{m}}_{r,t}^{D}=\alpha{\bm{m}}_{r,t-1}^{D}+\widehat{{\bm{g}}}^{D}_{r,t}=\sum_{i=0}^{t}\alpha^{t-i}\widehat{{\bm{g}}}^{D}_{r,i}\end{cases}$
@@
$\mathrm{(Shared~{}Momentum)}$$\begin{cases}-\theta^{S}_{f,t}=\sum_{i=0}^{t}{\bm{m}}^{S}_{f,i}+\sum_{i=0}^{t-1}{\bm{m}}^{S}_{r,i}\\ -\theta^{S}_{r,t}=\sum_{i=0}^{t}{\bm{m}}^{S}_{f,i}+\sum_{i=0}^{t}{\bm{m}}^{S}_{r,i}\end{cases}$ (11)
$\mathrm{(Decoupled~{}Momentum)}$$\begin{cases}-\theta^{D}_{f,t}=\sum_{i=0}^{t}{\bm{m}}^{D}_{f,i}+\sum_{i=0}^{t-1}{\bm{m}}^{D}_{r,i}\\ -\theta^{D}_{r,t}=\sum_{i=0}^{t}{\bm{m}}^{S}_{f,i}+\sum_{i=0}^{t}{\bm{m}}^{D}_{r,i}\\ \end{cases}$
@@
$\mathrm{(Shared~{}Momentum)}$$\begin{cases}-\theta_{f,t}^{S}=\sum_{i=0}^{t}A_{2(t-i)}\widehat{{\bm{g}}}^{S}_{f,i}+\sum_{i=0}^{t-1}A_{2(t-i)-1}\widehat{{\bm{g}}}^{S}_{r,i}\\ -\theta_{r,t}^{S}=\sum_{i=0}^{t}A_{2(t-i)+1}\widehat{{\bm{g}}}^{S}_{f,i}+\sum_{i=0}^{t}A_{2(t-i)}\widehat{{\bm{g}}}^{S}_{r,i}\\ \end{cases}$ (12)
$\mathrm{(Decoupled~{}Momentum)}$$\begin{cases}-\theta_{f,t}^{D}=\sum_{i=0}^{t}A_{t-i}\widehat{{\bm{g}}}^{D}_{f,i}+\sum_{i=0}^{t-1}A_{t-i-1}\widehat{{\bm{g}}}^{D}_{r,i}\\ -\theta_{r,t}^{D}=\sum_{i=0}^{t}A_{t-i}\widehat{{\bm{g}}}^{D}_{f,i}+\sum_{i=0}^{t}A_{t-i}\widehat{{\bm{g}}}^{D}_{r,i}\\ \end{cases}$
@@
[…]
@@
We assume negligible correlation between gradients from the forget loss and the retain loss in Assumption 3.2. Therefore, we can derive the variance for $\theta^{S}_{r,0}$ and $\theta^{D}_{r,0}$ as follows:
[…]
@@
That is to say, the conclusion of the theorem holds for $t=0$. Now we assume that the conclusion of the theorem holds for $10,1,...,t-$ 1, and then consider the case of $t$.
@@
$\forall w_{0},...,w_{t},$ $\mathrm{Var}\left(\sum_{i=0}^{t}w_{i}\widehat{{\bm{g}}}_{f,i}\right)=\sum_{i=0}^{t}w_{i}^{2}\mathrm{Var}(\widehat{{\bm{g}}}_{f,i})+2\sum_{i\neq j}\rho(\widehat{{\bm{g}}}_{f,i},\widehat{{\bm{g}}}_{f,j})w_{i}w_{j}\sqrt{\mathrm{Var}(\widehat{{\bm{g}}}_{f,i})\mathrm{Var}(\widehat{{\bm{g}}}_{f,j})}$ (17)

### v2

@@ 3.4 Dual Optimizers for Machine Unlearning
[…]
@@ 4.1 Random Subset Unlearning in Image Classification
Table 1: Performance summary of MU methods for image classification. Experiments are conducted on (a) $10\%$ random subset of CIFAR-10 using ResNet-18 and (b) $10\%$ random subset of TinyImageNet using Swin-T. All results are presented as mean and standard deviation across 55 trials with different random forget data. Performance gaps with RT are indicated in blue. The average gap (Gap) and average standard deviation (Std) metrics are calculated by the average of the gaps and standard deviation measured in FA, RA, TA, and MIA, respectively. All the numbers are in percentage.
@@
We start with random subset unlearning tasks in image classification, including using ResNet-18 (he2016deep, ) on CIFAR-10 (krizhevsky2009learning, ) and Swin Transformer-Tiny (Swin-T) (liu2021swin, ) on TinyImageNet (russakovsky2015imagenet, ). Additional results on CIFAR-100 (krizhevsky2009learning, ) and SVHN (netzer2011reading, ) are included in Appendix C.3. Consistent with previous work (thudi2022unrolling, ; jia2023model, ; fan2024salun, ; huang2025unified, ), we regard Retrain (RT) as the gold standard of MU. Since our proposed DualOptim is plug-and-play, we apply it to SCRUB (kurmanji2023towards, ), SalUn (fan2024salun, ), and SFRon (huang2025unified, ) to validate its efficacy and stability. Additionally, we include three simple baselines, namely Fine-tune (FT) (warnecke2021machine, ), Gradient Ascent (GA) (thudi2022unrolling, ) and Random Label (RL) (golatkar2020eternal, ), for reference as well. Note that the hyperparameters of all evaluated methods are tuned to their optimal values by sophisticated search, and we defer the implementation details to Appendix B. Besides the metrics mentioned in Sec. 3.1, which includes FA, RA, TA and MIA, we follow the evaluation criteria in (fan2024salun, ; huang2025unified, ) to report the average gap (Gap) and average standard deviation (Std) to indicate the overall performance and stability, respectively. They are the average gap with RT and the average standard deviation among the results in FA, RA, TA, and MIA, respectively.
@@ 4.2 Class-wise Unlearning in Image Generation
Table 2: Class-wise unlearning performance on CIFAR-10 with DDPM and ImageNet with DiT. The best unlearning performance for each forgetting class is highlighted in bold for FA (in %) and FID. Note that the results of SA, SalUn and SFRon are those reported in (huang2025unified, ).
Method CIFAR-10 Class-wise Unlearning
Automobile Cat Dog Horse Truck
FA $\downarrow$ FID $\downarrow$ FA $\downarrow$ FID $\downarrow$ FA $\downarrow$ FID $\downarrow$ FA $\downarrow$ FID $\downarrow$ FA $\downarrow$ FID $\downarrow$
SA 0.00 23.56 14.20 21.34 8.60 21.19 0.00 21.13 0.00 29.04
SalUn 0.20 21.23 1.40 20.29 0.00 20.18 0.60 20.70 0.80 20.45
SFRon 0.00 20.70 7.40 18.44 0.20 18.89 0.00 19.93 0.00 20.61
 +DO 0.20 19.72 1.00 19.36 0.00 18.58 0.00 18.91 0.00 17.26
Method ImageNet Class-wise Unlearning
Cockatoo Golden Retriever White Wolf Arctic Fox Otter
FA $\downarrow$ FID $\downarrow$ FA $\downarrow$ FID $\downarrow$ FA $\downarrow$ FID $\downarrow$ FA $\downarrow$ FID $\downarrow$ FA $\downarrow$ FID $\downarrow$
SA 0.00 348.75 0.00 298.97 0.00 45.89 0.00 393.91 29.8 321.21
SalUn 91.21 18.47 46.09 25.28 0.00 15.16 45.90 408.07 87.5 19.69
SFRon 0.00 13.59 0.00 17.76 0.00 23.28 0.00 16.12 0.00 16.43
 +DO 0.00 17.46 0.00 14.63 0.00 14.72 0.00 14.91 0.00 14.55
Besides classification, we further evaluate our proposed DualOptim in class-wise unlearning tasks in image generation. Following the settings of huang2025unified , we conduct experiments using conditional DDPM (ho2020denoising, ) with the U-Net (ronneberger2015u, ) on CIFAR-10 and the latent diffusion model (rombach2022high, ) with Diffusion Transformer (DiT) (peebles2023scalable, ) on ImageNet (russakovsky2015imagenet, ). We apply our proposed DualOptim to SFRon (huang2025unified, ), which exhibit the state-of-the-art unlearning performance in image generation. We also include SA (heng2023selective, ) and SalUn (fan2024salun, ) as baselines for comparison. Note that we do not apply DualOptim to SA and SalUn due to their joint updating scheme and unstable performance on ImageNet. Nonetheless, we still present the results of SalUn+DO on CIFAR-10 in Appendix C.4. The implementation details can be found in Appendix B. As for the evaluation metrics, we adopt the accuracy of the unlearned model’s generated images on forgetting classes (FA) by a pre-trained classifier to indicate the forgetting efficacy and Fréchet Inception Distance (FID) (heusel2017gans, ) to assess the generation capability on the retained classes.
@@ 4.3 Random Subset Unlearning in Large Language Models
Method Phi-1.5
forget 1% data forget 5% data forget 10% data
@@
GA+GD 0.4934 0.4493 0.4714 0.4360 0.5084 0.4722 0.4471 0.5246 0.4859
NPO+GD 0.2569 0.5682 0.4125 0.4940 0.4469 0.4705 0.4808 0.4382 0.4595
ME+GD 0.4944 0.3938 0.4441 0.4559 0.4480 0.4520 0.4594 0.4564 0.4579
 +DO 0.4866 0.6913 0.5889 0.4676 0.8200 0.6438 0.5009 0.7732 0.6370
DPO+GD 0.2410 0.6831 0.4621 0.4105 0.6334 0.5219 0.3517 0.6302 0.4910
IDK+AP 0.4403 0.5723 0.5063 0.4800 0.5112 0.4956 0.4614 0.6003 0.5308
 +DO 0.4221 0.7037 0.5629 0.4633 0.6974 0.5804 0.4422 0.7193 0.5807
Method LLaMA 2
forget 1% data forget 5% data forget 10% data
MC $\uparrow$ FE $\uparrow$ Avg. $\uparrow$ MC $\uparrow$ FE $\uparrow$ Avg. $\uparrow$ MC $\uparrow$ FE $\uparrow$ Avg. $\uparrow$
GA+GD 0.6696 0.5908 0.6302 0.0000 0.8772 0.4386 0.5592 0.9346 0.7469
NPO+GD 0.6414 0.6109 0.6262 0.5465 0.6921 0.6193 0.5648 0.7668 0.6658
ME+GD 0.7271 0.9204 0.8237 0.7472 0.9313 0.8392 0.7357 0.9489 0.8423
 +DO 0.7425 0.9612 0.8519 0.7316 0.9602 0.8459 0.7315 0.9625 0.8470
DPO+GD 0.7564 0.5335 0.6450 0.0000 0.8243 0.4122 0.0000 0.8041 0.4021
IDK+AP 0.7580 0.7625 0.7603 0.7529 0.7479 0.7504 0.7471 0.7433 0.7452
 +DO 0.7412 0.8075 0.7743 0.7354 0.7958 0.7656 0.7362 0.7855 0.7609
Following the success of MU in vision tasks, there is a rising interest of applying it to remove private or harmful information from large language models (LLMs). We conduct the experiments on Phi-1.5 (li2023textbooks, ) and LLaMA 2 (touvron2023llama, ), both are fine-tuned on TOFU dataset (mainitofu, ) with 1.3B and 7B parameters, respectively. We include three untargeted unlearning methods (GA+GD, NPO+GD and ME+GD) and two targeted unlearning methods (DPO+GD and IDK+AP) (yuan2024closer, ) as the baselines. We employ DualOptim (DO) in ME+GD and IDK+AP, which perform the best in (yuan2024closer, ), to validate the effectiveness of our method. Consistent with (mainitofu, ; yuan2024closer, ), we consider three levels of unlearning tasks: to forget 1%, 5%, and 10% of the constructed data. We follow (yuan2024closer, ) and adopt the improved Model Capability (MC)  [^2] Model Capability (MC) is referred to Model Utility (MU) in (yuan2024closer, ), which conflicts with the abbreviation of Machine Unlearning (MU) in this paper. and Forger Efficacy (FE) as the evaluation metrics.
@@
It is important to note that the forgetting and retaining objectives are jointly optimized in the baselines listed in Table 4, whereas our method alternates between optimizing these two objectives. To ensure a fair comparison, we also evaluate unlearning performance by alternately optimizing the forgetting and retaining objectives using a shared optimizer. The results in Table 15 of Appendix C.5 demonstrate that DualOptim surpasses both the joint and alternate update schemes.
Table 5: Running time and memory usage of Adam and DualOptim on Llama 2 with full-parameter tuning and LoRA. Note that for full-parameter tuning, we used two H20 GPUs and DeepSpeed zero stage 2 is adopted; for LoRA, we used a single H20 GPU.
Tuning method MU method Adam DualOptim Ratio
Full-parameter ME+GD 19.40 28.50 1.47
IDK+AP 22.77 33.92 1.49
LoRA ME+GD 17.04 17.43 1.02
IDK+AP 26.76 62.32 1.21
(a) Running Time Usage (min)
Tuning method MU method Adam DualOptim Ratio
Full-parameter IDK+AP 59.83 86.40 1.44
ME+GD 59.84 86.52 1.45
LoRA IDK+AP 30.63 30.62 1.00
ME+GD 31.44 30.62 0.97
(b) Memory Usage (GB/GPU)
Although DualOptim exhibits effectiveness in MU for LLMs, as shown in Table 5, the decoupled momentum introduces approximately $1.5\times$ computational and memory overhead because of large amount of parameters in LLMs. To investigate our method’s effectiveness under limited computation resources, we apply DualOptim to the parameter-efficient fine-tuning techniques, such as LoRA (hulora, ). As indicated in Table 16 of Appendix C.5, DualOptim achieves a minimal compromise in unlearning performance by using LoRA, while significantly reducing memory consumption on optimizer states.
@@ 4.4 Ablation Studies
We compare various combinations of the Adam and SGD optimizers for the forget loss and the retain loss with their standalone counterparts. Besides Adam, we also investigate the combinations of SGD and other optimizers, e.g., Lion (chen2023symbolic, ) and Muon (jordan2024muon, ). The results in Table 6 indicate that the configuration of Adam (F) + SGD (R) offers the optimal balance between performance and stability. This finding underscores the importance of decoupled momentum and adaptive learning rate for the forgetting objective. Moreover, the results presented in Table 18 and Table 19 in Appendix C.6 indicate that the best optimizer for retaining depends on the choice of optimizer during pretraining and the specific MU algorithms employed. It should be pointed out that a shared SGD demonstrates a competitive average performance but suffers from high standard deviation indicating instability. Conversely, Lion exhibits low variability but does not achieve optimal performance. These insights further emphasize the efficacy of our method in enhancing performance and stability in MU tasks.
Table 6: Ablation study on different optimizers when we use DualOptim in SFRon. Results are based on $10\%$ random subset unlearning task on CIFAR-10 using ResNet-18 pre-trained by SGD. (F) and (R) denotes that the optimizer is used to minimize the forget and retain losses, respectively.
@@ 5 Conclusion
Broader Impacts and Limitations
Our method contributes to the trustworthiness of deep learning and can be broadly applicable across diverse tasks. Although DualOptim can be integrated with parameter-efficient fine-tuning methods, it still introduces double memory consumption for optimizer states. We leave the development of a more efficient approach as future work.
Acknowledgments and Disclosure of Funding
This work is supported by National Natural Science Foundation of China (NSFC Project No. 62306250) and City University of Hong Kong (CityU Project No. 9220132).
@@ Appendix A Proofs
$\mathrm{Var}(\nabla_{\theta}\mathcal{L}(\theta))=\frac{1}{2}\iint p(\theta_{1})p(\theta_{2})\left\|\nabla\mathcal{L}(\theta_{1})-\nabla\mathcal{L}(\theta_{2})\right\|^{2}d\theta_{1}d\theta_{2}.$ (7)
@@
$\mathrm{Var}(\nabla_{\theta}\mathcal{L}(\theta))\leq\frac{L^{2}}{2}\iint p(\theta_{1})p(\theta_{2})\|\theta_{1}-\theta_{2}\|^{2}d\theta_{1}d\theta_{2}.$ (8)
Since $\mathrm{Var}(\theta)\leq\sigma_{\theta}^{2}$, we have $\frac{1}{2}\iint p(\theta_{1})p(\theta_{2})\|\theta_{1}-\theta_{2}\|^{2}d\theta_{1}d\theta_{2}\leq\sigma_{\theta}^{2}$. Therefore, we can conclude $\mathrm{Var}(\nabla_{\theta}{\mathcal{L}}(\theta))\leq L^{2}\sigma_{\theta}^{2}$
@@
[…]
@@
[…]
@@
[…]
@@
$\mathrm{Var}(\theta_{f,0}^{S})\leq\sigma^{2}\overset{\text{\tiny def}}{=}\overline{\mathrm{Var}}(\theta_{f,0}^{S}),\quad\mathrm{Var}(\theta_{f,0}^{D})\leq\sigma^{2}\overset{\text{\tiny def}}{=}\overline{\mathrm{Var}}(\theta_{f,0}^{D}).$ (13)
@@
We assume negative correlation between gradients from the forget loss and the retain loss in Assumption 3.2. Therefore, we can derive the variance for $\theta^{S}_{r,0}$ and $\theta^{D}_{r,0}$ as follows:
$\mathrm{Var}(\theta_{r,0}^{S})$ $\leq(1+\alpha)^{2}\mathrm{Var}(\widehat{{\bm{g}}}^{S}_{f,0})+\mathrm{Var}(\widehat{{\bm{g}}}^{S}_{r,0})\leq(1+\alpha)^{2}\sigma^{2}+(L^{2}+1)\sigma^{2}\overset{\text{\tiny def}}{=}\overline{\mathrm{Var}}(\theta^{S}_{r,0}),$ (15)
$\mathrm{Var}(\theta_{r,0}^{D})$ $\leq\mathrm{Var}(\widehat{{\bm{g}}}^{D}_{f,0})+\mathrm{Var}(\widehat{{\bm{g}}}^{D}_{r,0})\leq\sigma^{2}+(1+L^{2})\sigma^{2}\overset{\text{\tiny def}}{=}\overline{\mathrm{Var}}(\theta^{D}_{r,0}).$
@@
That is to say, the conclusion of the theorem holds for $t=0$. Now we assume that the conclusion of the theorem holds for 0,1,,t10,1,...,t-1, and then consider the case of $t$.
@@
$\forall w_{0},.,w_{t},$ $\mathrm{Var}\left(\sum_{i=0}^{t}w_{i}\widehat{{\bm{g}}}_{f,i}\right)=\sum_{i=0}^{t}w_{i}^{2}\mathrm{Var}(\widehat{{\bm{g}}}_{f,i})+2\sum_{i\neq j}\rho(\widehat{{\bm{g}}}_{f,i},\widehat{{\bm{g}}}_{f,j})w_{i}w_{j}\sqrt{\mathrm{Var}(\widehat{{\bm{g}}}_{f,i})\mathrm{Var}(\widehat{{\bm{g}}}_{f,j})}$ (17)
@@
A.3 Variance Bound of Performance Metric Function
We let $M(\theta)$ to represent the performance of model with parameter $\theta$ where $M$ can be FA/RA/TA/MIA. Before we derive the variance bound of performance metric function $M(\theta)$, we make the following assumption:
Assumption A.1.
The performance metric function $M$ is Lipschitz continuous:
$\forall\theta_{1},\theta_{2},\|M(\theta_{1})-M(\theta_{2})\|\leq L_{M}\|\theta_{1}-\theta_{2}\|,$ (21)
where $L_{M}$ is a non-negative constant.
Following a similar logic to Lemma 3.4, we have the following corollary:
Corollary A.2.
If Assumption A.1 holds, and the parameter variance satisfies $\mathrm{Var}(\theta)\leq\overline{\mathrm{Var}}(\theta)$, then we have $\mathrm{Var}(M(\theta))\leq L_{M}^{2}\overline{\mathrm{Var}}(\theta)$.
Proof.
Similar to the proof of Lemma 3.4, we have:
$\mathrm{Var}(M(\theta))=\frac{1}{2}\iint p(\theta_{1})p(\theta_{2})\left\|M(\theta_{1})-M(\theta_{2})\right\|^{2}d\theta_{1}d\theta_{2},$ (22)
where $\theta_{1},\theta_{2}$ are independent copies of $\theta$ with probability density $p(\theta)$.
By $\|M(\theta_{1})-M(\theta_{2})\|\leq L_{M}\|\theta_{1}-\theta_{2}\|$, we have:
$\mathrm{Var}(M(\theta))\leq\frac{L_{M}^{2}}{2}\iint p(\theta_{1})p(\theta_{2})\|\theta_{1}-\theta_{2}\|^{2}d\theta_{1}d\theta_{2}.$ (23)
Since $\mathrm{Var}(\theta)\leq\overline{\mathrm{Var}}(\theta)$, we have $\frac{1}{2}\iint p(\theta_{1})p(\theta_{2})\|\theta_{1}-\theta_{2}\|^{2}d\theta_{1}d\theta_{2}\leq\overline{\mathrm{Var}}(\theta)$. Therefore, we can conclude $\mathrm{Var}(M(\theta))\leq L_{M}^{2}\overline{\mathrm{Var}}(\theta)$. ∎
Discussion. The final unlearned parameters are usually the parameter updated by retaining loss at the last iteration $T$, i.e., $\theta^{D}_{r,T}$ using decoupled momentum and $\theta^{S}_{r,T}$ using shared momentum. According to Theorem 3.5, we have $\overline{\mathrm{Var}}(\theta^{D}_{r,T})\leq\overline{\mathrm{Var}}(\theta^{S}_{r,T})$. Let the upper bound of the performance variance $\overline{\mathrm{Var}}(M(\theta))=L_{M}^{2}\overline{\mathrm{Var}}(\theta)$, we can derive $\overline{\mathrm{Var}}(M(\theta^{D}_{r,T}))\leq\overline{\mathrm{Var}}(M(\theta^{S}_{r,T}))$. Thus, decoupled momentum can also induce less variability on performance metrics.